OpenStudy (anonymous):

sqrt(-98) * sqrt(-50) = ?

OpenStudy (anonymous):

\[\sqrt{-98}* \sqrt{-50}\\\sqrt{-2*49}* \sqrt{-2*25}\\7\sqrt{-2}* 5\sqrt{-2}\\35*-2\]

OpenStudy (anonymous):

ok, so an online tutor from college wrote it out like this...... sqrt(-98) * sqrt(-50)= sqrt(-1*98) * sqrt (-50)= a(p*q)=a(p) * a(a)= sqrt(-1) * sqrt(98) * sqrt(-1) * sqrt (50) sqrt(98)=i *sqrt (50)=i sqrt(49*2)=i *sqrt(25*2)=i and this is where i'm stuck